# Collision Avoidance and Geofencing for Fixed-wing Aircraft with Control Barrier Functions

### v2

@@ V Conclusion
In this paper, we developed a run-time assurance system for fixed-wing aircraft that intervenes into the operation of existing flight controllers whenever necessary for the safety of the aircraft. Specifically, we used control barrier functions (CBFs) to establish controllers with formal safety guarantees for collision avoidance, geofencing and the simultaneous execution thereof. We established and proved the safety guarantees provided by these controllers, and we demonstrated safe operation by numerical simulation of a nonlinear kinematic aircraft model and a high-fidelity dynamical model. We highlighted that different choices of CBFs—higher-order and backstepping-based CBFs that use model information, and simplified position-based CBF candidates in a model-free context—lead to qualitatively different behaviors while guaranteeing safety. As future work, we plan to enforce safety w.r.t. other constraints in flight envelopes such as angle of attack bounds.
@@ Appendix A Equations of Motion
Here we derive the equations of the 3D Dubins model (1)-(2). First, we revisit a six-degrees-of-freedom (6 DoF) model from [23] that governs the rigid-body dynamics of fixed-wing aircraft. Then, we simplify these dynamics to the 3D Dubins model through certain assumptions.
@@
$\begin{split}\mathbf{R}_{\rm eb}(\boldsymbol{\xi})&=\begin{bmatrix}{\rm c}\psi\!&\!-{\rm s}\psi\!&\!0\\ {\rm s}\psi\!&\!{\rm c}\psi\!&\!0\\ 0\!&\!0\!&\!1\end{bmatrix}\!\!\begin{bmatrix}{\rm c}\theta\!&\!0\!&\!{\rm s}\theta\\ 0\!&\!1\!&\!0\\ -{\rm s}\theta\!&\!0\!&\!{\rm c}\theta\end{bmatrix}\!\!\begin{bmatrix}1\!&\!0\!&\!0\\ 0\!&\!{\rm c}\phi\!&\!-{\rm s}\phi\\ 0\!&\!{\rm s}\phi\!&\!{\rm c}\phi\end{bmatrix}\!,\\ \mathbf{R}_{\rm bw}(\boldsymbol{\eta})&=\begin{bmatrix}{\rm c}\alpha\!&\!0\!&\!-{\rm s}\alpha\\ 0\!&\!1\!&\!0\\ {\rm s}\alpha\!&\!0\!&\!{\rm c}\alpha\end{bmatrix}\!\!\begin{bmatrix}{\rm c}\beta\!&\!-{\rm s}\beta\!&\!0\\ {\rm s}\beta\!&\!{\rm c}\beta\!&\!0\\ 0\!&\!0\!&\!1\end{bmatrix}\!,\\ \mathbf{R}_{\rm be}(\boldsymbol{\xi})&=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})^{\top},\quad\mathbf{R}_{\rm wb}(\boldsymbol{\eta})=\mathbf{R}_{\rm bw}(\boldsymbol{\eta})^{\top},\end{split}$ (57)
where ${\rm c}$ abbreviates $\cos$ and ${\rm s}$ abbreviates $\sin$.
@@ A-A Kinematics
$\mathbf{H}(\boldsymbol{\xi})=\begin{bmatrix}1&{\rm s}\phi{\rm t}\theta&{\rm c}\phi{\rm t}\theta\\ 0&{\rm c}\phi&-{\rm s}\phi\\ 0&{\rm s}\phi/{\rm c}\theta&{\rm c}\phi/{\rm c}\theta\end{bmatrix},$ (61)
and ${\rm t}$ abbreviates $\tan$. Moreover, after laborious calculation, it can be shown that the following useful identities hold:
$\begin{split}&\frac{\partial\mathbf{v}}{\partial\boldsymbol{\xi}}(\boldsymbol{\xi},\boldsymbol{\eta})\mathbf{H}(\boldsymbol{\xi})\boldsymbol{\omega}=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})\big{(}\boldsymbol{\omega}\times\mathbf{v}_{\rm b}(\boldsymbol{\eta})\big{)},\\ &\frac{\partial\mathbf{v}}{\partial\eta}(\boldsymbol{\xi},\boldsymbol{\eta})=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})\mathbf{R}_{\rm bw}(\boldsymbol{\eta})\begin{bmatrix}1&0&0\\ 0&V_{\rm T}&0\\ 0&0&V_{\rm T}{\rm c}\beta\end{bmatrix}.\end{split}$ (62)
@@
$\mathbf{v}(\boldsymbol{\xi},\boldsymbol{\eta})=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})\begin{bmatrix}V_{\rm T}\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}V_{\rm T}{\rm c}\theta{\rm c}\psi\\ V_{\rm T}{\rm c}\theta{\rm s}\psi\\ -V_{\rm T}{\rm s}\theta\end{bmatrix}.$ (64)
@@
$\mathbf{M}_{\mathbf{a}}(\boldsymbol{\xi},\boldsymbol{\eta})=\begin{bmatrix}\frac{\partial\mathbf{v}}{\partial V_{\rm T}}(\boldsymbol{\xi},\boldsymbol{\eta})\!&\!\begin{bmatrix}\frac{\partial\mathbf{v}}{\partial\theta}(\boldsymbol{\xi},\boldsymbol{\eta})\!&\!\frac{\partial\mathbf{v}}{\partial\psi}(\boldsymbol{\xi},\boldsymbol{\eta})\end{bmatrix}\mathbf{H}_{\theta\psi}(\boldsymbol{\xi})\end{bmatrix}\\ =\begin{bmatrix}{\rm c}\theta{\rm c}\psi\!\!&\!\!-V_{\rm T}({\rm c}\phi{\rm s}\theta{\rm c}\psi\!+\!{\rm s}\phi{\rm s}\psi)\!\!&\!\!V_{\rm T}({\rm s}\phi{\rm s}\theta{\rm c}\psi\!-\!{\rm c}\phi{\rm s}\psi)\\ {\rm c}\theta{\rm s}\psi\!\!&\!\!V_{\rm T}(-{\rm c}\phi{\rm s}\theta{\rm s}\psi\!+\!{\rm s}\phi{\rm c}\psi)\!\!&\!\!V_{\rm T}({\rm s}\phi{\rm s}\theta{\rm s}\psi\!+\!{\rm c}\phi{\rm c}\psi)\\ -{\rm s}\theta\!\!&\!\!-V_{\rm T}{\rm c}\phi{\rm c}\theta\!\!&\!\!V_{\rm T}{\rm s}\phi{\rm c}\theta\end{bmatrix}\!\!,$ (66)
@@ A-B Dynamics
[…]
@@ A-C Governing Equations
We simplify the 6 DoF model to the 3D Dubins model under the following assumptions: zero angle of attack, ${\alpha\equiv 0}$, zero angle of side slip, ${\beta\equiv 0}$, and zero crosswind force, ${C(\boldsymbol{\eta})\equiv 0}$. Furthermore, we neglect the dynamics of the angular velocity $\boldsymbol{\omega}$, while we use the longitudinal acceleration ${A_{\rm T}=\dot{V}_{\rm T}}$ directly instead of the thrust force $F_{\rm T}$. Then, the dynamics of $\beta$, $\alpha$, and $\boldsymbol{\omega}$ are dropped, leading to the 3D Dubins model:
@@
where ${\boldsymbol{\eta}=\begin{bmatrix}V_{\rm T}&0&0\end{bmatrix}^{\top}}$. By spelling out all terms, this gives (1). Furthermore, since ${\alpha\equiv 0}$, ${\beta\equiv 0}$ and ${C(\boldsymbol{\eta})\equiv 0}$, we obtain ${\mathbf{R}_{\rm wb}(\boldsymbol{\eta})=\mathbf{R}_{\rm bw}(\boldsymbol{\eta})=\mathbf{I}}$, and the second component of the body-axes force equations (69), ${\dot{\beta}\equiv 0}$, simplifies to (2). Moreover, the third component of (69), ${\dot{\alpha}\equiv 0}$, gives the required lift force, ${L=m(g_{\rm D}{\rm c}\phi{\rm c}\theta+V_{\rm T}Q)}$, that could keep ${\alpha\equiv 0}$. This shows that the assumption ${\alpha\equiv 0}$ implies a specific state-dependent lift force. Although this may not true in practice, we use this assumption as approximation to reduce the full dynamics model to simplified kinematics that represent the aircraft’s overall motion with less complexity.

### v3

@@ V Conclusion
In this paper, we developed a run-time assurance system for fixed-wing aircraft that intervenes into the operation of existing flight controllers whenever necessary for the safety of the aircraft. Specifically, we used control barrier functions (CBFs) to establish controllers with formal safety guarantees for collision avoidance, geofencing and the simultaneous execution thereof. We established and proved the safety guarantees provided by these controllers, and we demonstrated safe operation by numerical simulation of a nonlinear kinematic aircraft model and a high-fidelity dynamical model. We highlighted that different choices of CBFs—high-order and backstepping-based CBFs that use model information, and simplified position-based CBF candidates in a model-free context—lead to qualitatively different behaviors while guaranteeing safety. As future work, we plan to enforce safety w.r.t. other constraints in flight envelopes such as angle of attack bounds or attitude constraints. We also plan to formally address and enforce input constraints with techniques like the backup set method [15]. Finally, we plan to validate safety-critical controllers by hardware experiments.
@@ Appendix A Equations of Motion
Here we derive the equations of the 3D Dubins model (1)-(2). First, we revisit a six-degrees-of-freedom (6 DoF) model from [33] that governs the rigid-body dynamics of fixed-wing aircraft. Then, we simplify these dynamics to the 3D Dubins model through certain assumptions. While the upcoming equations are well-known in the literature, they are required by the controllers in the main body of the paper.
@@
$\begin{split}\mathbf{R}_{\rm eb}(\boldsymbol{\xi})&=\begin{bmatrix}{\rm c}_{\psi}\!&\!-{\rm s}_{\psi}\!&\!0\\ {\rm s}_{\psi}\!&\!{\rm c}_{\psi}\!&\!0\\ 0\!&\!0\!&\!1\end{bmatrix}\!\!\begin{bmatrix}{\rm c}_{\theta}\!&\!0\!&\!{\rm s}_{\theta}\\ 0\!&\!1\!&\!0\\ -{\rm s}_{\theta}\!&\!0\!&\!{\rm c}_{\theta}\end{bmatrix}\!\!\begin{bmatrix}1\!&\!0\!&\!0\\ 0\!&\!{\rm c}_{\phi}\!&\!-{\rm s}_{\phi}\\ 0\!&\!{\rm s}_{\phi}\!&\!{\rm c}_{\phi}\end{bmatrix}\!,\\ \mathbf{R}_{\rm bw}(\boldsymbol{\eta})&=\begin{bmatrix}{\rm c}_{\alpha}\!&\!0\!&\!-{\rm s}_{\alpha}\\ 0\!&\!1\!&\!0\\ {\rm s}_{\alpha}\!&\!0\!&\!{\rm c}_{\alpha}\end{bmatrix}\!\!\begin{bmatrix}{\rm c}_{\beta}\!&\!-{\rm s}_{\beta}\!&\!0\\ {\rm s}_{\beta}\!&\!{\rm c}_{\beta}\!&\!0\\ 0\!&\!0\!&\!1\end{bmatrix}\!,\\ \mathbf{R}_{\rm be}(\boldsymbol{\xi})&=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})^{\top},\quad\mathbf{R}_{\rm wb}(\boldsymbol{\eta})=\mathbf{R}_{\rm bw}(\boldsymbol{\eta})^{\top},\end{split}$ (55)
where ${\rm c}_{(.)}$ abbreviates $\cos(.)$ and ${\rm s}_{(.)}$ abbreviates $\sin(.)$.
@@ A-A Kinematics
$\mathbf{H}(\boldsymbol{\xi})=\begin{bmatrix}1&{\rm s}_{\phi}{\rm t}_{\theta}&{\rm c}_{\phi}{\rm t}_{\theta}\\ 0&{\rm c}_{\phi}&-{\rm s}_{\phi}\\ 0&{\rm s}_{\phi}/{\rm c}_{\theta}&{\rm c}_{\phi}/{\rm c}_{\theta}\end{bmatrix},$ (59)
and ${\rm t}_{\theta}$ abbreviates $\tan(\theta)$. Moreover, after laborious calculation, it can be shown that the following useful identities hold:
$\begin{split}&\frac{\partial\mathbf{v}}{\partial\boldsymbol{\xi}}(\boldsymbol{\xi},\boldsymbol{\eta})\mathbf{H}(\boldsymbol{\xi})\boldsymbol{\omega}=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})\big{(}\boldsymbol{\omega}\times\mathbf{v}_{\rm b}(\boldsymbol{\eta})\big{)},\\ &\frac{\partial\mathbf{v}}{\partial\eta}(\boldsymbol{\xi},\boldsymbol{\eta})=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})\mathbf{R}_{\rm bw}(\boldsymbol{\eta})\begin{bmatrix}1&0&0\\ 0&V_{\rm T}&0\\ 0&0&V_{\rm T}{\rm c}_{\beta}\end{bmatrix}.\end{split}$ (60)
@@
$\mathbf{v}(\boldsymbol{\xi},\boldsymbol{\eta})=\mathbf{R}_{\rm eb}(\boldsymbol{\xi})\begin{bmatrix}V_{\rm T}\\ 0\\ 0\end{bmatrix}=\begin{bmatrix}V_{\rm T}{\rm c}_{\theta}{\rm c}_{\psi}\\ V_{\rm T}{\rm c}_{\theta}{\rm s}_{\psi}\\ -V_{\rm T}{\rm s}_{\theta}\end{bmatrix}.$ (62)
@@
$\mathbf{M}_{\mathbf{a}}(\boldsymbol{\xi},\boldsymbol{\eta})=\begin{bmatrix}\frac{\partial\mathbf{v}}{\partial V_{\rm T}}(\boldsymbol{\xi},\boldsymbol{\eta})\!&\!\begin{bmatrix}\frac{\partial\mathbf{v}}{\partial\theta}(\boldsymbol{\xi},\boldsymbol{\eta})\!&\!\frac{\partial\mathbf{v}}{\partial\psi}(\boldsymbol{\xi},\boldsymbol{\eta})\end{bmatrix}\mathbf{H}_{\theta\psi}(\boldsymbol{\xi})\end{bmatrix}\\ =\begin{bmatrix}{\rm c}_{\theta}{\rm c}_{\psi}\!\!&\!\!-V_{\rm T}({\rm c}_{\phi}{\rm s}_{\theta}{\rm c}_{\psi}\!+\!{\rm s}_{\phi}{\rm s}_{\psi})\!\!&\!\!V_{\rm T}({\rm s}_{\phi}{\rm s}_{\theta}{\rm c}_{\psi}\!-\!{\rm c}_{\phi}{\rm s}_{\psi})\\ {\rm c}_{\theta}{\rm s}_{\psi}\!\!&\!\!V_{\rm T}(-{\rm c}_{\phi}{\rm s}_{\theta}{\rm s}_{\psi}\!+\!{\rm s}_{\phi}{\rm c}_{\psi})\!\!&\!\!V_{\rm T}({\rm s}_{\phi}{\rm s}_{\theta}{\rm s}_{\psi}\!+\!{\rm c}_{\phi}{\rm c}_{\psi})\\ -{\rm s}_{\theta}\!\!&\!\!-V_{\rm T}{\rm c}_{\phi}{\rm c}_{\theta}\!\!&\!\!V_{\rm T}{\rm s}_{\phi}{\rm c}_{\theta}\end{bmatrix}\!\!,$ (64)
@@ A-B Dynamics
$\begin{split}&\mathbf{f}_{\boldsymbol{\eta}}(\boldsymbol{\xi},\boldsymbol{\eta},\boldsymbol{\omega})=\mathbf{A}(\boldsymbol{\eta})\Big{(}\!-\!\frac{1}{m}\mathbf{F}_{\rm A}(\boldsymbol{\eta})+\mathbf{R}_{\rm wb}(\boldsymbol{\eta})\mathbf{R}_{\rm be}(\boldsymbol{\xi})\mathbf{g}_{\rm D}\\ &\hskip 85.35826pt+\mathbf{v}_{\rm w}(\boldsymbol{\eta})\times\big{(}\mathbf{R}_{\rm wb}(\boldsymbol{\eta})\boldsymbol{\omega}\big{)}\Big{)},\\ &\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{\eta})\!=\!\mathbf{A}(\boldsymbol{\eta})\frac{1}{m}\mathbf{R}_{\rm wb}(\boldsymbol{\eta})\!\!\begin{bmatrix}1\\ 0\\ 0\end{bmatrix}\!\!,\quad\mathbf{A}(\boldsymbol{\eta})\!=\!\begin{bmatrix}1\!\!&\!\!0\!\!&\!\!0\\ 0\!\!&\!\!\frac{1}{V_{\rm T}}\!\!&\!\!0\\ 0\!\!&\!\!0\!\!&\!\!\frac{1}{V_{\rm T}{\rm c}_{\beta}}\end{bmatrix}\!\!.\end{split}$ (68)
@@ A-C Governing Equations
We simplify the 6 DoF model to the 3D Dubins model using the following assumptions.
Assumption 3.
The angle of attack, the angle of side slip, and the crosswind force are zero: ${\alpha\equiv 0}$, ${\beta\equiv 0}$, and ${C(\boldsymbol{\eta})\equiv 0}$. The dynamics of the angular velocity $\boldsymbol{\omega}$ are neglected.
Furthermore, we use the longitudinal acceleration ${A_{\rm T}=\dot{V}_{\rm T}}$ directly instead of the thrust force $F_{\rm T}$. Then, the dynamics of $\beta$, $\alpha$, and $\boldsymbol{\omega}$ are dropped, leading to the 3D Dubins model:
@@
where ${\boldsymbol{\eta}=\begin{bmatrix}V_{\rm T}&0&0\end{bmatrix}^{\top}}$. By spelling out all terms, this gives (1). Furthermore, since ${\alpha\equiv 0}$, ${\beta\equiv 0}$, and ${C(\boldsymbol{\eta})\equiv 0}$, we obtain ${\mathbf{R}_{\rm wb}(\boldsymbol{\eta})=\mathbf{R}_{\rm bw}(\boldsymbol{\eta})=\mathbf{I}}$, and the second component of the body-axes force equations (67), ${\dot{\beta}\equiv 0}$, simplifies to (2). Moreover, the third component of (67), ${\dot{\alpha}\equiv 0}$, gives the required lift force, ${L=m(g_{\rm D}{\rm c}_{\phi}{\rm c}_{\theta}+V_{\rm T}Q)}$, that could keep ${\alpha\equiv 0}$. This shows that the assumption ${\alpha\equiv 0}$ implies a specific state-dependent lift force. Although this may not true in practice, we use this assumption as approximation to reduce the full dynamics model to simplified kinematics that represent the aircraft’s overall motion with less complexity.